a tesla car needs to charge its 350-volt model and must send about 715 amps through its cabling to the battery. how many electrons pass through the cable for the current involved.


1
Expert's answer
2021-05-05T15:48:15-0400

Current is defined as "I=\dfrac{\Delta Q}{\Delta t}"


So we can find the total charge: 

"\Delta Q= Current\times time\ period"

Lets say car is charged for 1 minutes

So, Total charge transferred in 1 minutes= "715\times 60 =42900\ C"

and number of electrons pass through wire in one minutes="\dfrac{42900\ C}{1.6\times 10^{-19}\frac{C}{electrons}}=2.68\times 10^{23}\ \ electrons"